# SoulX-LiveTalk: Real-Time Infinite Streaming of Audio-Driven Avatars via Self-Correcting Bidirectional Distillation

### v2

@@ Abstract
Deploying massive diffusion models for real-time, infinite-duration, audio-driven avatar generation presents a significant engineering challenge, primarily due to the conflict between computational load and strict latency constraints. Existing approaches often compromise visual fidelity by enforcing strictly unidirectional attention mechanisms or reducing model capacity. To address this problem, we introduce SoulX-LiveTalk, a 14B-parameter framework optimized for high-fidelity real-time streaming. Diverging from conventional unidirectional paradigms, we use a Self-correcting Bidirectional Distillation strategy that retains bidirectional attention within video chunks. This design preserves critical spatiotemporal correlations, significantly enhancing motion coherence and visual detail. To ensure stability during infinite generation, we incorporate a Multi-step Retrospective Self-Correction Mechanism, enabling the model to autonomously recover from accumulated errors and preventing collapse. Furthermore, we engineered a full-stack inference acceleration suite incorporating hybrid sequence parallelism, Parallel VAE, and kernel-level optimizations. Extensive evaluations confirm that SoulX-LiveTalk is the first 14B-scale system to achieve a sub-second start-up latency (0.87s) while reaching a real-time throughput of 32 FPS, setting a new standard for high-fidelity interactive digital human synthesis.
@@ 1 Introduction
To address these challenges, we introduce SoulX-LiveTalk, a low-latency, real-time, audio-driven avatar framework built upon a 14B-parameter DiT. Compared with prior approaches, our method offers substantial advantages across model architecture design, training and distillation strategy, and inference-time optimization, enabling high-fidelity avatar generation while meeting the stringent latency requirements of live streaming.
@@
Second, in terms of training efficiency, SoulX-LiveTalk streamlines the time-consuming training paradigm of LiveAvatar (huang2025live). This improvement is enabled by the architectural consistency described above, which allows us to adopt a lightweight yet effective training strategy. Latency-Aware Spatiotemporal Adaptation adapts the model to operate effectively under lower spatial resolutions and shorter temporal horizons for real-time constraints. Self-Correcting Bidirectional Distillation further reduces inference steps and eliminates classifier-free guidance, while introducing a multi-step retrospective self-correction mechanism inspired by Self-Forcing++ to enable the model to recover from self-induced deviations and prevents collapse during long-horizon generation.
@@ 2 SoulX-LiveTalk
This section details the core methodology of SoulX-LiveTalk. As illustrated in Figure 2, our framework is built upon a 14B-parameter DiT and integrates a two-stage training pipeline with a full-stack inference acceleration engine. The training protocol progresses from Latency-Aware Spatiotemporal Adaptation phase to a Self-Correcting Bidirectional Distillation phase, designed to satisfy both high-fidelity generation and low-latency streaming requirements.
@@ 2.2 Model Training
Following this two-stage training strategy, SoulX-LiveTalk achieves state-of-the-art performance in both inference speed and generation quality compared to existing audio-driven video generation models.
@@ 3 Experiment
3.1 Performance of SoulX-LiveTalk
SoulX-LiveTalk against state-of-the-art audio-driven generation models, including Ditto (li2025ditto), EchoMimic-V3 (meng2025echomimicv3), StableAvatar (tu2025stableavatar), OmniAvatar (gan2025omniavatar), InfiniteTalk (yang2025infinitetalk), and LiveAvatar (huang2025live).
@@ 3.1.1 Quantitative Analysis
Table 1: Quantitative comparison of SoulX-LiveTalk and state-of-the-art methods on TalkBench-Short ($10$ s) and TalkBench-Long (>5>5 min). Models marked with support real-time inference, and denotes the LoRA version.
@@
SoulX-LiveTalk (Ours) 3.51 4.79 1.47 11.56 99.22 98.10 99.61 99.52 32.00
@@
SoulX-LiveTalk (Ours) 3.12 3.04 1.61 12.25 99.29 98.36 99.63 99.58 32.00
Table. 1 compares SoulX-LiveTalk with state-of-the-art methods on the TalkBench-Short and TalkBench-Long datasets.
On the short-video benchmark, SoulX-LiveTalk achieves the highest scores in visual quality and synchronization. Specifically, it records an ASE of $3.51$ and an IQA of $4.79$, surpassing the previous best performer, Echomimic-V3, which scored $3.45$ and $4.70$ respectively. In terms of lip-sync precision, our model attains a Sync-C score of $1.47$, outperforming the $1.32$ score of OmniAvatar. For inference speed, the system achieves a throughput of $32$ FPS on a 14B-parameter model. This exceeds the real-time requirement of $25$ FPS and significantly outperforms the $20.88$ FPS recorded by LiveAvatar.
Regarding temporal consistency metrics such as Subject-C and BG-C, Ditto records the highest values across both datasets, including a Subject-C of $99.80$. This performance is attributed to Ditto’s generation paradigm which inpaints only the facial region while keeping the background and torso pixel-wise static. While this approach maximizes stability scores, it precludes the generation of full-body dynamics. SoulX-LiveTalk is designed to synthesize audio-driven full-body motion which naturally introduces greater pixel variance. Despite this increased complexity, it maintains a Subject-C score of $99.22$, demonstrating a balance between motion expressiveness and temporal stability.
For long-form generation, we assess robustness through synchronization retention. SoulX-LiveTalk achieves a Sync-C of $1.61$ and a Sync-D of $12.25$. These scores are better than InfiniteTalk and LiveAvatar. Additionally, the model sustains a throughput of $32$ FPS in long-duration tasks. These results confirm that the bidirectional distillation strategy effectively reduces the desynchronization and drift often found in unidirectional streaming models.
@@ 3.1.2 Qualitative Analysis
This section presents a qualitative assessment of SoulX-LiveTalk, focusing on generation fidelity, long-term stability, and lip-sync precision.
@@
Figure 3: Visual quality comparison on 5-second video generation. Orange boxes highlight static hand poses in Ditto, while blue boxes highlight significant artifacts (e.g., hand distortion, over-exposure) in baselines. In contrast, SoulX-LiveTalk eliminates these artifacts, demonstrating superior structural integrity and detail fidelity.
@@
Visual Fidelity and Detail Preservation. Figure. 3 compares the visual quality of 5-second video generations. Baseline models demonstrate significant struggles with plausible dynamics during large limb movements. Ditto fails to synthesize meaningful hand motion, with poses remaining static throughout the sequence, as highlighted in orange. Echomimic-v3 and StableAvatar exhibit structural distortions and artifacts in hand regions, marked in blue. InfiniteTalk suffers from hand over-exposure and excessive motion blur during rapid gestures. In contrast, SoulX-LiveTalk leverages its 14B DiT architecture and bidirectional attention mechanism to eliminate these artifacts. It synthesizes clear, structurally sound hand movements with sharp textures, avoiding the issues observed in baselines. Furthermore, our method surpasses LiveAvatar in background consistency and identity fidelity.
Stability in Infinite Generation.
@@
Figure 4: Qualitative evaluation of long-term stability. Comparison across $10$s to $1000$s reveals structural collapse in baselines (blue boxes) versus the sustained robustness of SoulX-LiveTalk (orange boxes), which preserves sharp details even after 1000 seconds of continuous generation.
@@
Figure. 4 evaluates generation stability over continuous sequences extending to 1000 seconds. Baseline models, including LiveAvatar, StableAvatar, and InfiniteTalk, suffer from significant error accumulation over time. As indicated by the blue boxes, these methods exhibit severe texture blurring and loss of detail in background regions. SoulX-LiveTalk mitigates error propagation through bidirectional streaming distillation and its self-correction mechanism. As shown in the orange boxes, our model preserves consistent facial geometry and sharp background details at the 1000-second mark, verifying its robustness for infinite streaming.
@@
Fine-grained Lip-sync Precision. Figure. 5 assesses lip-sync fidelity during specific Chinese phonetic articulations. Baseline methods struggle with complex phonemes, exhibiting structural misalignment. As highlighted by the yellow dashed boxes, during the articulation of characters such as “上” (shàng) and “突” (tū), competitors fail to match the mouth aperture and shape of the Ground Truth (GT), resulting in visible distortions. Conversely, SoulX-LiveTalk captures these fine-grained phonemic dynamics, yielding lip geometries that are rigorously aligned with the GT. This precision minimizes lip-sync drift and stiffness, ensuring visual authenticity across different languages.
@@ 4 Conclusion and Future Work
We presented SoulX-LiveTalk, a framework designed to meet real-time requirements while maintaining high-quality video synthesis. Integrating Bidirectional Streaming Distillation with a Multi-step Self-Correction mechanism allows our 14B-parameter DiT model to sustain stable, infinite-length streaming on an $\timesH8$$0$ cluster. Our approach also simplifies training. We demonstrate that complex multi-stage pre-training is not required. A brief SFT phase followed by distribution matching distillation suffices to achieve state-of-the-art performance. We are open-sourcing this solution to serve as a practical baseline for the community.

### v3

# SoulX-FlashTalk: Real-Time Infinite Streaming of Audio-Driven Avatars via Self-Correcting Bidirectional Distillation

@@ Abstract
Deploying massive diffusion models for real-time, infinite-duration, audio-driven avatar generation presents a significant engineering challenge, primarily due to the conflict between computational load and strict latency constraints. Existing approaches often compromise visual fidelity by enforcing strictly unidirectional attention mechanisms or reducing model capacity. To address this problem, we introduce SoulX-FlashTalk, a 14B-parameter framework optimized for high-fidelity real-time streaming. Diverging from conventional unidirectional paradigms, we use a Self-correcting Bidirectional Distillation strategy that retains bidirectional attention within video chunks. This design preserves critical spatiotemporal correlations, significantly enhancing motion coherence and visual detail. To ensure stability during infinite generation, we incorporate a Multi-step Retrospective Self-Correction Mechanism, enabling the model to autonomously recover from accumulated errors and preventing collapse. Furthermore, we engineered a full-stack inference acceleration suite incorporating hybrid sequence parallelism, Parallel VAE, and kernel-level optimizations. Extensive evaluations confirm that SoulX-FlashTalk is the first 14B-scale system to achieve a sub-second start-up latency (0.87s) while reaching a real-time throughput of 32 FPS, setting a new standard for high-fidelity interactive digital human synthesis.
1 Introduction
@@
To address these challenges, we introduce SoulX-FlashTalk, a low-latency, real-time, audio-driven avatar framework built upon a 14B-parameter DiT. Compared with prior approaches, our method offers substantial advantages across model architecture design, training and distillation strategy, and inference-time optimization, enabling high-fidelity avatar generation while meeting the stringent latency requirements of live streaming.
@@
Second, in terms of training efficiency, SoulX-FlashTalk streamlines the time-consuming training paradigm of LiveAvatar (huang2025live). This improvement is enabled by the architectural consistency described above, which allows us to adopt a lightweight yet effective training strategy. Latency-Aware Spatiotemporal Adaptation adapts the model to operate effectively under lower spatial resolutions and shorter temporal horizons for real-time constraints. Self-Correcting Bidirectional Distillation further reduces inference steps and eliminates classifier-free guidance, while introducing a multi-step retrospective self-correction mechanism inspired by Self-Forcing++ to enable the model to recover from self-induced deviations and prevents collapse during long-horizon generation.
@@ 2 SoulX-FlashTalk
This section details the core methodology of SoulX-FlashTalk. As illustrated in Figure 2, our framework is built upon a 14B-parameter DiT and integrates a two-stage training pipeline with a full-stack inference acceleration engine. The training protocol progresses from Latency-Aware Spatiotemporal Adaptation phase to a Self-Correcting Bidirectional Distillation phase, designed to satisfy both high-fidelity generation and low-latency streaming requirements.
@@ 2.2 Model Training
Following this two-stage training strategy, SoulX-FlashTalk achieves state-of-the-art performance in both inference speed and generation quality compared to existing audio-driven video generation models.
@@ 3 Experiment
3.1 Performance of SoulX-FlashTalk
SoulX-FlashTalk against state-of-the-art audio-driven generation models, including Ditto (li2025ditto), EchoMimic-V3 (meng2025echomimicv3), StableAvatar (tu2025stableavatar), OmniAvatar (gan2025omniavatar), InfiniteTalk (yang2025infinitetalk), and LiveAvatar (huang2025live).
@@ 3.1.1 Quantitative Analysis
Table 1: Quantitative comparison of SoulX-FlashTalk and state-of-the-art methods on TalkBench-Short ($10$ s) and TalkBench-Long (>5>5 min). Models marked with support real-time inference, and denotes the LoRA version.
@@
SoulX-FlashTalk (Ours) 3.51 4.79 1.47 11.56 99.22 98.10 99.61 99.52 32.00
@@
SoulX-FlashTalk (Ours) 3.12 3.04 1.61 12.25 99.29 98.36 99.63 99.58 32.00
Table. 1 compares SoulX-FlashTalk with state-of-the-art methods on the TalkBench-Short and TalkBench-Long datasets.
On the short-video benchmark, SoulX-FlashTalk achieves the highest scores in visual quality and synchronization. Specifically, it records an ASE of $3.51$ and an IQA of $4.79$, surpassing the previous best performer, Echomimic-V3, which scored $3.45$ and $4.70$ respectively. In terms of lip-sync precision, our model attains a Sync-C score of $1.47$, outperforming the $1.32$ score of OmniAvatar. For inference speed, the system achieves a throughput of $32$ FPS on a 14B-parameter model. This exceeds the real-time requirement of $25$ FPS and significantly outperforms the $20.88$ FPS recorded by LiveAvatar.
Regarding temporal consistency metrics such as Subject-C and BG-C, Ditto records the highest values across both datasets, including a Subject-C of $99.80$. This performance is attributed to Ditto’s generation paradigm which inpaints only the facial region while keeping the background and torso pixel-wise static. While this approach maximizes stability scores, it precludes the generation of full-body dynamics. SoulX-FlashTalk is designed to synthesize audio-driven full-body motion which naturally introduces greater pixel variance. Despite this increased complexity, it maintains a Subject-C score of $99.22$, demonstrating a balance between motion expressiveness and temporal stability.
For long-form generation, we assess robustness through synchronization retention. SoulX-FlashTalk achieves a Sync-C of $1.61$ and a Sync-D of $12.25$. These scores are better than InfiniteTalk and LiveAvatar. Additionally, the model sustains a throughput of $32$ FPS in long-duration tasks. These results confirm that the bidirectional distillation strategy effectively reduces the desynchronization and drift often found in unidirectional streaming models.
@@ 3.1.2 Qualitative Analysis
This section presents a qualitative assessment of SoulX-FlashTalk, focusing on generation fidelity, long-term stability, and lip-sync precision.
@@
Figure 3: Visual quality comparison on 5-second video generation. Orange boxes highlight static hand poses in Ditto, while blue boxes highlight significant artifacts (e.g., hand distortion, over-exposure) in baselines. In contrast, SoulX-FlashTalk eliminates these artifacts, demonstrating superior structural integrity and detail fidelity.
@@
Visual Fidelity and Detail Preservation. Figure. 3 compares the visual quality of 5-second video generations. Baseline models demonstrate significant struggles with plausible dynamics during large limb movements. Ditto fails to synthesize meaningful hand motion, with poses remaining static throughout the sequence, as highlighted in orange. Echomimic-v3 and StableAvatar exhibit structural distortions and artifacts in hand regions, marked in blue. InfiniteTalk suffers from hand over-exposure and excessive motion blur during rapid gestures. In contrast, SoulX-FlashTalk leverages its 14B DiT architecture and bidirectional attention mechanism to eliminate these artifacts. It synthesizes clear, structurally sound hand movements with sharp textures, avoiding the issues observed in baselines. Furthermore, our method surpasses LiveAvatar in background consistency and identity fidelity.
@@
Figure 4: Qualitative evaluation of long-term stability. Comparison across $10$s to $1000$s reveals structural collapse in baselines (blue boxes) versus the sustained robustness of SoulX-FlashTalk (orange boxes), which preserves sharp details even after 1000 seconds of continuous generation.
@@
Stability in Infinite Generation. Figure. 4 evaluates generation stability over continuous sequences extending to 1000 seconds. Baseline models, including LiveAvatar, StableAvatar, and InfiniteTalk, suffer from significant error accumulation over time. As indicated by the blue boxes, these methods exhibit severe texture blurring and loss of detail in background regions. SoulX-FlashTalk mitigates error propagation through bidirectional streaming distillation and its self-correction mechanism. As shown in the orange boxes, our model preserves consistent facial geometry and sharp background details at the 1000-second mark, verifying its robustness for infinite streaming.
@@
Fine-grained Lip-sync Precision. Figure. 5 assesses lip-sync fidelity during specific Chinese phonetic articulations. Baseline methods struggle with complex phonemes, exhibiting structural misalignment. As highlighted by the yellow dashed boxes, during the articulation of characters such as “上” (shàng) and “突” (tū), competitors fail to match the mouth aperture and shape of the Ground Truth (GT), resulting in visible distortions. Conversely, SoulX-FlashTalk captures these fine-grained phonemic dynamics, yielding lip geometries that are rigorously aligned with the GT. This precision minimizes lip-sync drift and stiffness, ensuring visual authenticity across different languages.
@@ 4 Conclusion and Future Work
We presented SoulX-FlashTalk, a framework designed to meet real-time requirements while maintaining high-quality video synthesis. Integrating Bidirectional Streaming Distillation with a Multi-step Self-Correction mechanism allows our 14B-parameter DiT model to sustain stable, infinite-length streaming on an $\timesH8$$0$ cluster. Our approach also simplifies training. We demonstrate that complex multi-stage pre-training is not required. A brief SFT phase followed by distribution matching distillation suffices to achieve state-of-the-art performance. We are open-sourcing this solution to serve as a practical baseline for the community.